OpenStudy (vera_ewing):

Solve 81^x = 27^(x + 2) x = 1 x = 2 x = 5 x = 6

OpenStudy (jdoe0001):

\(\large \bf { 81^x = 27^{x+2}\qquad {\color{brown}{ 3^4=81\qquad 27=3^3}}\qquad (3^4)^x=(3^3)^{x+2} \\ \quad \\ 3^{4\cdot x}=3^{3\cdot (x+2)} }\)